3. Calculate the price of a $1,000 10 year 5% coupon bond that yields 6%.MUST SHOW ALL WORK!!
Answer #1

Face Value(FV) = $1000

Time Period = 10 years

Coupon (C) = 5% = $50

Yield to Maturity (Y) = 6%

Price of Bond = Present Value of Future Cash Flows

=Present value of ( Coupon Payment + Face Value on Maturity)

= Σ(c/(1 + Υ):) + FV/(1 + Υ) 10

=c (1/(1+y)) + FV/(1+ y ) 16

= 50*(1/1.06)(1-(1/1.06)^10)/(1-(1/1.06) + 1000/(1.06)^10 [ based of sum of finite GP i.e. a*(1 –p)/(1-r)

   with a = 1/1.06 and r=1/1.06

= (50/1.06)*(0.441605)/0.056604 + 1000/1.79085

= 47.17*7.802 + 558.395

=368.005 + 558.395

= 926.4

Thus Bond price is $926.40